# Electron charge coherence on a solid neon surface

### v1

@@ IV.1 Phonon-induced displacement of neon surface
Here $K_{1}(\eta)$ is the modified Bessel function of the first kind. The first two terms describe the kinematic interaction arising from the electron’s wavefunction being set to zero on a non-flat hard-wall surface. The first two terms have nonzero contribution only if we are considering out-of-plane excitation. The pressing electric field term in practical experiments is much smaller than the polarization term. So, we only need to include the polarization term in our calculation.
The relaxation rate can be calculated with Eq. (27),
@@
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (37)
and
@@
$\mathit{\Gamma}^{(\text{dis})}$ $=\frac{R^{2}r_{\scriptscriptstyle\text{B}}^{2}\omega_{0}^{6}}{8\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{9}}\int_{0}^{1}{\rm d}\gamma\,\gamma^{2}(1-\gamma^{2})^{3}$
×$\times\exp\left[-\frac{\hbar\omega_{0}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
×$\times\left[\ln\left(\frac{\omega_{0}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}\sqrt{1-\gamma^{2}}\right)\right]^{2}.$ (39)
For simplicity, we have assumed again $\omega_{x}=\omega_{y}=\omega_{0}$. Numerical integral shows $\mathit{\Gamma}^{(\text{dis})}=49.4$ s-1 for the transition frequency at $f_{0}=6.4$ GHz. This leads to the corresponding relaxation time,
$T_{1}^{\left(\text{dis}\right)}=1/\mathit{\Gamma}^{\left(\text{dis}\right)}=20.2\text{ ms},$ (40)
and the coherence time,
$T_{2}^{\left(\text{dis}\right)}=\left(\frac{\mathit{\Gamma}^{\left(\text{dis}\right)}}{2}+\Gamma_{\phi}\right)^{-1}=40.4\text{ ms},$ (41)
@@
Moreover, as seen from Eq. (39), the decay rate $\mathit{\Gamma}^{\left(\text{dis}\right)}$ is dependent on $\omega_{x}$. In Table IV.1 and Fig.2, we present the variations of $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ with the transition frequency $f_{x}$. In most cases, $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ exceed 1 ms, which surpasses the $T_{1}$ and $T_{2}$ values for most semiconductor as well as superconducting charge qubits.
@@ IV.2 Phonon-induced modulation of dielectric constant
Another decoherence mechanism comes from the phonon-induced modulation of dielectric constant ^$\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$. The dielectric constant variation $\delta\epsilon$ and the density variation $\varrho$ is related by
@@
×$\times K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|),$ (43)
which is independent of $\bm{r}_{\scriptscriptstyle\parallel}$ and contains the modified Bessel function of the first kind, $K_{1}(\eta)$.
The relaxation rate can be calculated with Eq. (27),
@@
×$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}).$ (44)
@@
×$\left.\times\int_{-\infty}^{0}{\rm d}z^{\prime}\,\frac{\sin(q_{z}z^{\prime})}{z-z^{\prime}}K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|)\right|^{2},$ (45)
where $\psi_{1}(z)$ is the normalized electron’s ground-state wavefunction in $z$ given by Eq. (6). To calculate the integrals in $F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$, since $z>0$ for the electron and $z^{\prime}<0$ for the image charges always hold, $z-z^{\prime}=|z-z^{\prime}|>0$ always holds.
@@
$\mathit{\Gamma}^{\text{(mod)}}$ $=\frac{8R^{2}\omega_{0}^{4}}{\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{7}}\int_{0}^{1}{\rm d}\gamma\,(1-\gamma^{2})$
×$\times\exp\left[-\frac{\hbar\omega_{0}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
×$\times\left|\int_{0}^{\infty}{\rm d}s\int_{0}^{\infty}{\rm d}s^{\prime}\,\frac{s^{2}}{(s+s^{\prime})^{2}}\sin\left(\frac{\omega_{0}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}us^{\prime}\right){\rm e}^{-2s}\right|^{2},\,$ (46)
where $\gamma\equiv\cos\vartheta$ with $\vartheta$ being the polar angle of phonon wavevectors, $s\equiv z/r_{\scriptscriptstyle\text{B}}>0$, and $s^{\prime}\equiv-z^{\prime}/r_{\scriptscriptstyle\text{B}}>0$. They are all positive dimensionless variables. For simplicity, we have assumed $\omega_{x}=\omega_{y}=\omega_{0}$. Eq. (46) can be numerically integrated without further approximations.
For a qubit transition frequency of $f_{0}=\omega_{0}/2\pi=6.4$ GHz, as commonly used in our experiments, ignoring the practical need of anharmonicity, the numerical integration gives a relaxation rate of $\mathit{\Gamma}^{\text{(mod)}}=908$ s-1. Consequently, the relaxation time $T_{1}$ is
$T_{1}^{\text{(mod)}}=\frac{1}{\mathit{\Gamma}^{\text{(mod)}}}=1.1\,\text{ms}.$ (47)
Hence, the coherence time $T_{2}$ is relaxation-limited by phonon emission [35],
$T_{2}^{\text{(mod)}}=\left(\frac{\mathit{\Gamma}^{\text{(mod)}}}{2}+\mathit{\Gamma}_{\phi}\right)^{-1}=2.2\,\text{ms}.$ (48)
These calculated numbers are at least one order of magnitude better than our experimentally observed $\sim$0.1 ms. Moreover, the decay rate $\mathit{\Gamma}^{(\text{mod})}$ has a complicated dependence on $\omega_{0}$, not only in the prefactor but also in the integrals. In Table 2 and Fig.2 , we present how $T_{1}$ and $T_{2}$ vary with the qubit transition frequency $f_{0}$. In all cases, they are longer than 0.1 ms, and longer than those of most semiconductor charge qubits.
Table 2. : Calculated relaxation time $T_{1}^{\text{(mod)}}$ and coherence time $T_{2}^{\text{(mod)}}$ of an electron in-plane motional state with varied transition frequency $f_{0}$ for phonon-induced dielectric-constant modulation.
$f_{0}$ (GHz) $T_{1}^{\text{(mod)}}$ (s) $T_{2}^{\text{(mod)}}(s)$
1 13.4913.4913.49 27.027.027.0
2 0.330.330.33 0.660.660.66
3 0.0410.0410.041 0.0820.0820.082
4 0.0100.0100.010 $1.98\times 10^{-2}$
@@
So far, we have assumed that the solid Ne occupies the entire lower half space $z<0$. This is a valid approximation when the Ne film is thicker than $\sim$10 nm. However, if the Ne film is only a few nanometers thick, the image charge will be dominated by the underlying dielectric substrate, such as a silicon (Si) or sapphire. Then the electron decoherence is primarily induced by the phonons in Si or sapphire. At low temperatures, the longitudinal sound velocity in silicon and sapphire are $c_{\text{Si}}=8480$ m/s [36] and $c_{\text{Sa}}=11350$ m/s, respectively [37]. Consequently, the typical electron’s in-plane wavenumber $\left(m\omega_{x}/\hbar\right)^{1/2}=1.9\times 10^{7}$ m-1 is much larger than the wavenumbers of the resonant phonons $\omega_{0}/c_{\text{Si}}=4.7\times 10^{6}$ m-1 and $\omega_{0}/c_{\text{Sa}}=3.5\times 10^{6}$ m-1 ($f_{0}=6.4$ GHz). This mismatch makes the relaxation rate through these channels essentially 0, following the above calculation. Physically it originates from an inability to simultaneously satisfy energy and momentum conservation during the scattering process. Therefore, the single phonon scattering from the substrate is exponentially suppressed due to the mismatch between the size of the electron wave function and the phonon wavelength at the same energy. In experiments, the thickness of the Ne film typically lies between the two aforementioned extremes. As a result, the actual electron charge coherence time may exceed our conservative estimations above.
@@ V Conclusion
In our comprehensive theoretical study, we examined the decoherence mechanisms impacting the motional states of a single electron on a solid Ne surface at a temperature of 101010 mK. Notably, we found that the relaxation time for electron motional states with a transition frequency of 6.4 GHz is $T_{1}^{\left(\text{vac}\right)}=\,99$ s in vacuum and $T_{1}^{\left(\text{cav}\right)}=\,32$ ms in a superconducting cavity with coupling constant $g\sim 2\pi\times 5$ MHz, resonator linewidth $\kappa\sim 2\pi\times 0.5$ MHz, and detuning from resonant frequency $\left|f_{x}-f_{\text{r}}\right|\sim 500$ MHz. The coherence time $T_{2}^{\left(\text{vac}\right)}$ has upper bounds of 198 s and 64 ms, respectively. Additionally, we investigated the decoherence processes caused by phonon-induced displacement of the Ne surface, finding coherence times $T_{2}^{\left(\text{dis}\right)}$ ranging from approximately 366 s to 7 ms, spanning the transition frequency from 1 GHz to 10 GHz. Furthermore, we discovered that phonon-induced modulations of Ne relative permittivity can significantly alter the coherence time $T_{2}^{\left(\text{mod}\right)}$ from 27 s to 0.3 ms, within the same transiton frequency range. These results highlight the promising potential of single-electron motional (charge) qubits on solid Ne that may outperform most existing semiconductor and superconducting charge qubits.

### v2

@@ IV.1 Phonon-induced displacement of neon surface
Here $K_{1}(\eta)$ is the modified Bessel function of the first kind. The first two terms describe the kinematic interaction arising from the electron’s wavefunction being set to zero on a non-flat hard-wall surface. The first two terms have nonzero contribution only if we are considering out-of-plane excitation. The pressing electric field term in practical experiments is much smaller than the polarization term, so only the polarization term needs to be included in our calculation. The relaxation rate can be simplified as
@@
$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}),$ (35)
where
@@
$\mathit{\Gamma}^{(\text{dis})}$ $=\frac{R^{2}r_{\scriptscriptstyle\text{B}}^{2}\omega_{x}^{6}}{8\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{9}}\int_{0}^{1}{\rm d}\gamma\,\gamma^{2}(1-\gamma^{2})^{3}$
$\times\exp\left[-\frac{\hbar\omega_{x}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
$\times\left[\ln\left(\frac{\omega_{x}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}\sqrt{1-\gamma^{2}}\right)\right]^{2}.$ (37)
For simplicity, we have assumed $\omega_{x}=\omega_{y}$. The decay rate $\mathit{\Gamma}^{\left(\text{dis}\right)}$ is solely determined by the confinement frequency $\omega_{x}$. Numerical integral shows $\mathit{\Gamma}^{(\text{dis})}=49.4$ s-1 for the transition frequency at $f_{x}=6.4$ GHz. This leads to the corresponding relaxation time,
$T_{1}^{\left(\text{dis}\right)}(6.4\,\text{GHz})=1/\mathit{\Gamma}^{\left(\text{dis}\right)}=20.2\text{ ms},$ (38)
and coherence time,
$T_{2}^{\left(\text{dis}\right)}(6.4\,\text{GHz})=\left(\frac{\mathit{\Gamma}^{\left(\text{dis}\right)}}{2}+\Gamma_{\phi}\right)^{-1}=40.4\text{ ms},$ (39)
@@
In Table IV.1 and Fig.2, we present the variations of $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ with different transition frequency $f_{x}$. In most cases, $T_{1}^{\left(\text{dis}\right)}$ and $T_{2}^{\left(\text{dis}\right)}$ exceed 1 ms, which surpass the $T_{1}$ and $T_{2}$ values for most semiconductor and superconducting charge qubits.
@@ IV.2 Phonon-induced modulation of dielectric constant
Another decoherence mechanism comes from the phonon-induced modulation of dielectric constant $\hat{\mathcal{H}}_{\text{el-ph}}^{(\text{mod})}$. The dielectric constant variation $\delta\epsilon$ and the density variation $\varrho$ is related by
@@
$\times K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|),$ (41)
which is independent of $\bm{r}_{\scriptscriptstyle\parallel}$ and contains the modified Bessel function of the first kind, $K_{1}(\eta)$. The relaxation rate can be calculated with Eq. (26),
@@
$\times\delta(\hbar\omega_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}-\hbar\omega_{x}).$ (42)
@@
$\left.\times\int_{-\infty}^{0}{\rm d}z^{\prime}\,\frac{\sin(q_{z}z^{\prime})}{z-z^{\prime}}K_{1}(q_{\scriptscriptstyle\parallel}|z-z^{\prime}|)\right|^{2},$ (43)
where $\psi_{1}(z)$ is the normalized electron’s ground-state wavefunction in $z$ given by Eq. (6). To calculate the integrals in $F_{\bm{q}_{\scriptscriptstyle\parallel},q_{z}}$, note that $z>0$ for the electron and $z^{\prime}<0$ for the image charges, therefore $z-z^{\prime}=|z-z^{\prime}|>0$ always holds.
@@
$\mathit{\Gamma}^{\text{(mod)}}$ $=\frac{8R^{2}\omega_{x}^{4}}{\pi m_{e}\rho_{\scriptscriptstyle\text{Ne}}c_{\scriptscriptstyle\text{Ne}}^{7}}\int_{0}^{1}{\rm d}\gamma\,(1-\gamma^{2})$
$\times\exp\left[-\frac{\hbar\omega_{x}}{2m_{e}c_{\scriptscriptstyle\text{Ne}}^{2}}(1-\gamma^{2})\right]$
$\times\left|\int_{0}^{\infty}{\rm d}s\int_{0}^{\infty}{\rm d}s^{\prime}\,\frac{s^{2}}{(s+s^{\prime})^{2}}\sin\left(\frac{\omega_{x}}{c_{\scriptscriptstyle\text{Ne}}}r_{\scriptscriptstyle\text{B}}us^{\prime}\right){\rm e}^{-2s}\right|^{2},\,$ (44)
where $\gamma\equiv\cos\vartheta$ with $\vartheta$ being the polar angle of phonon wavevectors, $s\equiv z/r_{\scriptscriptstyle\text{B}}>0$, and $s^{\prime}\equiv-z^{\prime}/r_{\scriptscriptstyle\text{B}}>0$. They are all positive dimensionless variables and Eq. (44) can be numerically integrated without further approximations.
For a qubit transition frequency of $f_{x}=\omega_{x}/2\pi=6.4$ GHz, as commonly used in our experiments, ignoring the practical need of anharmonicity, the numerical integration gives a relaxation rate of $\mathit{\Gamma}^{\text{(mod)}}=908$ s-1. Consequently, the relaxation time $T_{1}$ is
$T_{1}^{\text{(mod)}}(6.4\,\text{GHz})=\frac{1}{\mathit{\Gamma}^{\text{(mod)}}}=1.1\,\text{ms}.$ (45)
and the coherence time $T_{2}$ is relaxation-limited [37],
$T_{2}^{\text{(mod)}}(6.4\,\text{GHz})=\left(\frac{\mathit{\Gamma}^{\text{(mod)}}}{2}+\mathit{\Gamma}_{\phi}\right)^{-1}=2.2\,\text{ms}.$ (46)
These calculated numbers are at least one order of magnitude better than our experimentally observed $\sim$0.1 ms. In Table IV.2 and Fig. IV.2, we present how $T_{1}$ and $T_{2}$ vary with the qubit transition frequency $f_{x}$. In all cases, they are longer than 0.1 ms, and longer than those of most semiconductor charge qubits.
Table 2. : Calculated relaxation time $T_{1}^{\text{(mod)}}$ and coherence time $T_{2}^{\text{(mod)}}$ of an electron in-plane motional state with varied transition frequency $f_{x}$ for phonon-induced dielectric-constant modulation.
$f_{x}$ (GHz) $T_{1}^{\text{(mod)}}$ (s) $T_{2}^{\text{(mod)}}(s)$
1 $13.49$ $27.0$
2 $0.33$ $0.66$
3 $0.041$ $0.082$
4 $0.010$ $1.98\times 10^{-2}$
@@
So far, we have assumed that the solid Ne occupies the entire lower half space $z<0$. This is a valid approximation when the Ne film is thicker than $\sim$10 nm. However, if the Ne film is only a few nanometers thick, the image charge will be dominated by the underlying dielectric substrate, such as a silicon (Si) or sapphire. Then the electron decoherence is primarily induced by the phonons in Si or sapphire. At low temperatures, the longitudinal sound velocity in silicon and sapphire are $c_{\text{Si}}=8480$ m/s (ref. [38]) and $c_{\text{Sa}}=11350$ m/s (ref. [39]), respectively. Consequently, the typical electron’s in-plane wavenumber $\left(m\omega_{x}/\hbar\right)^{1/2}=1.9\times 10^{7}$ m-1 is much larger than the wavenumbers of the resonant phonons $\omega_{x}/c_{\text{Si}}=4.7\times 10^{6}$ m-1 and $\omega_{x}/c_{\text{Sa}}=3.5\times 10^{6}$ m-1 ($f_{x}=6.4$ GHz). This mismatch makes the relaxation rate through these channels essentially 0, following the above calculation. Physically it originates from an inability to simultaneously satisfy energy and momentum conservation during the scattering process. Therefore, the single phonon scattering from the substrate is exponentially suppressed due to the mismatch between the size of the electron wave function and the phonon wavelength at the same energy. In experiments, the thickness of the Ne film typically lies between the two aforementioned extremes. As a result, the actual electron charge coherence time may exceed our conservative estimations above.
@@ V Conclusion
In this theoretical study, we have investigated the charge decoherence of a single electron on a solid neon surface interacting with the acoustic phonons in the solid neon. We find in the transition frequency range from 1 GHz to 10 GHz, for the phonon-induced displacement of the Ne surface, the coherence time $T_{2}^{\left(\text{dis}\right)}$ range from approximately 366 s to 7 ms, and for the phonon-induced modulations of Ne relative permittivity, the coherence tim $T_{2}^{\left(\text{mod}\right)}$ ranges from 27 s to 0.3 ms. These results highlight the promise of single-electron motional (charge) qubits on solid Ne as a long-coherence platform, calling for efforts to improve the experimentally achieved coherence performance.